OpenStudy (anonymous):

solve for x x+5 1 ----- = --- 6 3

OpenStudy (anonymous):

Cross multiply then solve for x: \[3(x+5)=6(1)\] \[3x+15=6\] \[x=-3\]

OpenStudy (anonymous):

x+5 = (1*6)/3 x+5 = 6/3 x+5 = 2 x = 2-5 x = -3